OpenStudy (anonymous):

Balance these equations: C5H10+O2....CO2+H2O KCIO3.....KCI+O2 NaOH+CO2...Na2CO3+H2O Cu2O+H2....Cu+H2O Fe+H2O.....Fe3O4+H2 CH3OH+O2......CO2+H2O AgNO3+BaCl 2.....Ba(NO3)2+AgCl H3PO4+NanOH.....Na3PO4+H2O P4+O2..P4O10 Zn+HCI...ZnCl2+H2 SiO2+C.....SiC+CO2 MgO+H2SO4.....MgSO4+H2O As4+O2.....As4O6 Cu+AgNO3.....Cu(NO3)2+Ag C2H5OH+O2....CO2+H2O H2S+O2.....H2O+SO2 Al2O3+HCI..AlCl3+AlCl2+H2O MnO2....Mn3O4+O2 Al+Fe2O3....Al2O3+Fe Na+H2O...NaOH+H2

OpenStudy (anonymous):

I already got the answers for the first 9, but i have no idea if they are correct or not.

OpenStudy (zpupster):

post your work and then someone might comment

OpenStudy (anonymous):

ok, one moment, here is the first 9

OpenStudy (anonymous):

Zn + 2HCl..ZnCl2+H2 SiO2+C....SiC+CO2 (balanced) MgO+H2SO4...MgSO4+H2O (balanced) As4+O2......As4O6 (balanced? because 4+2=6) 2Cu+AgNO3......Cu(NO3)2+Ag

OpenStudy (zpupster):

As4+O2......As4O6 (balanced? because 4+2=6)????? As bal your O's are not

OpenStudy (anonymous):

C2 H5 O2 H+O2.....CO2+H2O ? H2S+O2.......H2O+SO2 (balanced) Al2 O3+H2 CI......AlCl3+H2O

OpenStudy (anonymous):

As4+O2........As4 2O6 ?????

OpenStudy (zpupster):

no try again

OpenStudy (zpupster):

hint change left side

OpenStudy (zpupster):

2Cu+AgNO3......Cu(NO3)2+Ag this needs work (NO3)2 one on left side 2 on right??

OpenStudy (anonymous):

As4+O2........As4 2O6 ? but how can i change the 6? it makes no sense, since there are only 2 O molecules?

OpenStudy (zpupster):

As4+3O2........As4 2O6 3*O2 = 6 O's

OpenStudy (anonymous):

so: As4+3O2......As4O6

OpenStudy (zpupster):

you have to multiply the subscript by coefficient

OpenStudy (anonymous):

ok, so O is now correct, since 2*3=6, and the right side of the equation has 6 O molecules

OpenStudy (zpupster):

yes

OpenStudy (anonymous):

ok, thanks!

OpenStudy (zpupster):

2Cu+AgNO3......Cu(NO3)2+Ag this needs work (NO3)2 one on left side 2 on right?? do this because i have to go

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Cu+AgNO3......Cu(NO3)2+Ag Is it already balanced???

OpenStudy (zpupster):

no you have (NO3)2 on the right and only 1 on the left

OpenStudy (anonymous):

so: Cu+2AgNO3....Cu(NO3)2+Ag

OpenStudy (zpupster):

now you have 2 Ag's on right and one on left

OpenStudy (anonymous):

Cu+Ag 2NO3........Cu(NO3)2+Ag or would I need "( )"???

OpenStudy (zpupster):

Cu+2AgNO3......Cu(NO3)2+2Ag that is balance 1Cu L--1Cu rt 2Ag L--2Ag rt 2NO3 L --2NO3 Rt gotta run good luck!! just count

OpenStudy (abb0t):

Omg! That's...a lot of questions. lOL

OpenStudy (adamconner):

which ones do you still need help with?